CH 12 Information: Use the following to complete the CH 12 tab of the excel workbook provided on Blackboard.

Browning Underwriting and Recapitalization Professionals (BURP) buys and sells securities in short-term periods. The following transactions relate to BURP activity in December of 2019, which has a fiscal year end of December 31. All of the following securities meet the criteria to be classified as trading securities and no other trading securities were held by BURP prior to December of 2019.    

Dec 4      Purchased LRB Corp Bonds for $8 million

Dec 7      Purchased 2 million shares of Cupcake International for $14 million

Dec 10    Sold the LRB Corp Bonds for $8.5 million

Dec 15    Purchased US Treasury Bills for $47 million and US Treasury Bonds for $85 million

Dec 17    Sold 1.5 million shares of Cupcake International for $12 million

Dec 21    Sold the US Treasury Bills for $43 million

Dec 23    Sold the US Treasury Bonds for $82 million

Dec 27    Received payment of cash dividends from Cupcake International common shares for $250,000

Dec 31    Recorded any necessary adjusting/closing entries relating to the investments. Market price of Cupcake International was $10 per share.

  1. Prepare all appropriate journal entries for December 2019 in the excel template on the CH 12 tab.
  2. Calculate the total to be reported on the 2019 Balance Sheet for short-term investments.
  3. Calculate the total impact to profit reported on the 2019 Income Statement that is a result of the investment activity (i.e., the impact to profit for the investment activity).

1) Record all journal entries by day for December of 2019

2) Calculate the total to be reported on the 2019 Balance Sheet for short-term investments

Account title

Dr

Cr

Dec. 4

Dec. 7th

3) Calculate the total impact to profit reported on the 2019 Income Statement as a result of the investment activity
Answer #1


JOURNAL ENTRY Credit LRB CORP BONOS Decy / $8m Debit Date Account Detail $ 8 m cash Dect. I Share Of Cupcare international $1(LOSING / ADJUSTING PENTRIES Decio | Profit on sale or Bonds 1$.5ml PROCIT & LOBAL $.5m Declz $1.5m I profit on sale of Shane

2. Balance to be reported on the 2019 Balance Sheet for short-term investment:

* LRB Bonds are fully disposed, so, it'll not appear on balance sheet

* Us Treasury Bill & Bond are also disposed, so, it'll also not appear on balance sheet

NOTE-Investment, here being a current asset, should be valued at cost price or market price, whichever is lower. In this case Cost Price was $7($14 million/2 million), but the market price was $ 10. Hence, $7 should be taken into consideration

Remaining share of cupcake international= Purchased share-sold share

=2 million-1.5 million=.5 million or 500,000 shares

cost price of 500,000 share=500,000*7=3,500,000

so, short term investment will be shown at $3,500,000 on balance sheet.

3. Impact of profit on income statement

income from dividend=250,000

Profit on sale of bonds=$ .5 million or $ 500,000

profit on sale of shares of cupcake= $ 1.5 million or 1,500,000

total profit= 500,000+1,500,000=2,250,000

Less

Loss on sale of us treasury bill= $ 4 million or 4,000,000

Loss on sale of us treasury bond= $ 3 million or 3,000,000

total loss=4,000,000+3,000,000=7,000,000

net effect on income statement=Profit - loss

2,250,000-7,000,000=(-4,750,000) (LOSS)